OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line that contains the point Q( 1, -2) and is parallel to the line whose equation is y - 4 = 2/3 (x - 3)

OpenStudy (anonymous):

\[y - 4 = \frac{ 2 }{ 3 }x - 2\]\[y - \frac{ 2 }{ 3 }x - 2 =0\] \[\frac{ 2 }{ 3 }x - y + 2 =0\] \[2x - 3y + 6 = 0\]

OpenStudy (anonymous):

Nw Find the Solpe of This Line - a/b ie m1 = 2/3

OpenStudy (anonymous):

Using Point Slope Form \[ y - 1 = 2/3 (x +2)\] \[y = \frac{ 2 }{3 }x + \frac{ 4 }{ 3 } + 1\]

OpenStudy (anonymous):

Note : Since the Lines Are Parallel They Will have Same Slope